Ravi the trainer has two solo workout plans that he offers his clients: Plan A and Plan B. Each client does either one or the other (not both). On Friday there were 4clients who did Plan A and 8who did Plan B. On Saturday there were 2clients who did Plan A and 3
who did Plan B. Ravi trained his Friday clients for a total of 9hours and his Saturday clients for a total of 4hours. How long does each of the workout plans last?

Answer by mila nedic(19)

Let's say number of hours of plan A is represented by a and number of hours of plan B is represented by b. Let's represent Friday's and Saturday's hours using the variable a and b with equations.
Friday: 4a+8b=9
Saturday: 2a+3b=4
2a+3b=4
2a=4-3b (we make one variable on the each side)
4a=8-6b (now we can replace 4a with 8-6b in the Friday equation so we are only dealing with one variable
4a+8b=9
(8-6b)+8b=9
8+2b=9
2b=1
b=0.5 (plan b is 30 mins or 0.5 hours)
Now we can use b=0.5 to solve for a.
2a+3b=4
2a+3(0.5)=4
2a+1.5=4
2a=2.5
a=1.25 (plan a is 75 mins of 1.25 hours)
There fore Plan A is 75 minutes and Plan B is 30 minutes.